Say we have 14 people in this room; it was found there are 4 people having a height from
150 to 160 cm; 4 people having a height from 160 to 170 cm; 6 people having a height from 170 to 180 cm; what is the estimated height average in the class?
1
Expert's answer
2020-04-22T18:35:12-0400

Let "X" be a random variable such that "X(p)=\text{height of $p$}" for every pupil "p" in the room.

We have "P(X\in[150,160])=\frac{4}{14}=\frac{2}{7}", "P(X\in[160,170])=\frac{4}{14}=\frac{2}{7}", "P(X\in[170,180])=\frac{6}{14}=\frac{3}{7}"

Alos suppose that there is an uniform distribution on "[150,160]", "[160,170]" and "[170,180]".

So "\rho_X(x)=\begin{cases}
0,&\text{if $x\not\in[150,180]$}\\
\frac{1}{35},&\text{if $x\in[150,160)$}\\
\frac{1}{35},&\text{if $x\in[160,170)$}\\
\frac{3}{70},&\text{if $x\in[170,180]$}
\end{cases}="

"=\begin{cases}
0,&\text{if $x\not\in[150,180]$}\\
\frac{1}{35},&\text{if $x\in[150,170)$}\\
\frac{3}{70},&\text{if $x\in[170,180]$}
\end{cases}", where "\rho_X" is probability density function of "X".

So average height is "EX=\int\limits_{\mathbb R}x \rho_X(x)dx=\int\limits_{150}^{170}\frac{1}{35}xdx+\int\limits_{170}^{180}\frac{3}{70}xdx="

"=\frac{1}{70}x^2\bigl|_{150}^{170}+\frac{3}{140}x^2\bigl|_{170}^{180}=\frac{1165}{7}"

Answer: "\frac{1165}{7}" cm